OpenStudy (anonymous):

Marie has $830 in her bank account and withdraws $60 each month. Denise has $970 in her bank account and withdraws $80 each month. In how many months will Marie and Denise have equal amounts of money in their accounts? A. 4 B. 5 C. 7 D. 9

OpenStudy (anonymous):

Solve the following for m, the number of months to effect equality. 830 - 60 m = 970 - 80 m m = 7

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

Your welcome.